. Mr. Ali has Rs. 35000 income per month, the expenditure on pulses is 2000, when income
rises to 40000, the expenditure on pulses reduces to 1500 rupees, categorize the pulses as
inferior, superior or normal good.
1
Expert's answer
2020-07-07T12:17:00-0400

The pulses are inferior good to Mr. Ali.

Am inferior good is a good whose consumption decreases as the level of income increases. Based on the question, the level of consumption of pulses is decreasing with an increase in the level of income.

The demand of both normal and superior goods increases with an increase in the level of income.

Lets mathematically prove our answer above by calculating the income elasticity of demand.

Using the arc elasticity:

"\text{Income elasticity of demand}=\dfrac{\dfrac{Q2-Q1}{\dfrac{Q1+Q2}{2}}}{\dfrac{I2-I1}{\dfrac{I1+I2}{2}}}"


"\text{Income elasticity of demand}=\dfrac{\dfrac{1500-2000}{\dfrac{2000+1500}{2}}}{\dfrac{40000-35000}{\dfrac{35000+40000}{2}}}"


"\text{Income elasticity of demand}=\blue{-2.143}"

An inferior good is always associated with a negative elasticity of demand while a normal and a superior good is always associated with a positive income elasticity of demand. Therefore, the good in the question is an inferior good since it has a negative income elasticity of demand.